# Holstein mechanism in single-site model with unitary evolution

### v3

@@ Abstract
We investigate the Holstein mechanism in a single-site model, where unitary evolution intrinsically involves both fermion and boson operators under nonadiabatic conditions. The resulting unitary dynamics and boson-frequency dependence reveal a quantum phase transition, evidenced by distinct short-time (power-law decay) and long-time (exponential decay) behaviors, which are manifested in the polaronic shift, bosonic energy, and dynamics of reduced density matrix. This observation is consistent with a non-Markovian to Markovian transition. Also, in terms of the density operator, the additional perturbation from the degenerated ground state of the total Hamiltonian would cause fluctuations that outweigh the Markovian dynamics at long times. This may lead to non-unitary evolution and highly mixed states due to entanglement with the environment.
@@ I Introduction
Adiabaticity and non-adiabaticity fundamentally influence the physical properties of quantum systems, including their stability and phase transitions [1]. In the adiabatic limit, where the electron’s Fermi velocity significantly exceeds the sound velocity, electron-phonon scattering is elastic (even in the neutral limit), and interband transitions vanish, preventing electron energy loss. However, in the nonadiabatic regime, when a non-relativistic electron moves faster than the sound wave while still interacting with acoustic phonons, it can lose energy via Cherenkov radiation. For weak electron-phonon coupling in the stable (adiabatic) regime, the dispersion modification is linear with the impurity momentum. Conversely, for strong electron-phonon coupling, non-adiabaticity leads to an avoided crossing in the polaron band structure. Phenomena dependent on impurity motion, such as coherence or decoherence, are crucial for understanding polaron formation, as they are intrinsically linked to Fermi liquid or non-Fermi liquid behavior. For example, avoided crossings also appear in ultracold Fermi atomic systems under strong-interaction limits [6, 7, 8, 9]. Decoherence can occur in extreme cases, like a very light impurity immersed in a bath of heavy particles. Conversely, for a heavy impurity in a 1D system, decoherence can also arise at zero temperature due to the orthogonality catastrophe [2, 3, 4]. In the adiabatic case, the presence of phonon absorption and emission similarly induces decoherence, reducing coherent band motion.
In this paper, we study the single-site Holstein model. We initially choose the boson number operator with degenerated spectrum, which means the states do not distinguished by the particle number (population) but the internal degrees of freedom. When the perturbation weakly breaks the degeneracy, purity of the corresponding density matrix increases. where the internal degrees of freedom plays the main role. Also, by virtue of the high purity of the reduced density matrices, we gain the convenience of not needing to consider the deviation from the minimum uncertainty condition in the phase space.
Specifically, a time-dependent perturbation breaks the degeneracies of the original composite Hilbert spaces (electron and boson Hilbert spaces), causing the unitary evolution of the total Hamiltonian. This perturbation not only introduces time-dependence but also a Hermitian part to the boson operators. This Hermitian part of the boson operators, in turn, not only causes a nonzero commutation between the electron and boson terms but also leads the ground state of the annihilation boson operator to deviate from the classical coherent Gaussian state. The Markovian process for the system (non-adiabaticity) dominates, especially in the long-time limit. Here, the system’s relaxation time scale (characterized by exponential decay) is much slower than that of the bath electrons. Conversely, at short times, a finite non-Markovian effect with power-law decay is numerically proven, despite the absence of fluctuations as a result of non-local correlations due to the "single-site" restriction.
@@ II Model
We consider the model described by
$H=H_{e}+H_{b}+H_{eb},$ (1)
where $H_{e}(t)=\epsilon c^{{\dagger}}(t)c(t)$ is the single electron term. Due to the single-site consideration, the bare bandwidth is irrelevant here. Note that ${\rm Tr}[(c^{{\dagger}}(t)c(t))^{2}]={\rm Tr}[c^{{\dagger}}(t)c(t)]^{2}/N$, precluding the Hubbrd-type interacion effect. We consider the total Hamiltonian in a N-by-N Hilbert space, and a time-dependent perturbation cause breaks the degeneracies of original composite Hilbert spaces. This guarantees the unitary evolution of $H(t)$, which is distinct from the Heisenberg picture. The bosonic term reads
$H_{b}(t)=\frac{p^{2}(t)}{2}+\frac{1}{2}\omega^{2}x^{2}(t)=\omega(b^{{\dagger}}(t)b(t)+\frac{{\bf I}}{2}),$ (2)
where
$b(t)=\sqrt{\frac{\omega}{2}}(x(t)+\frac{ip(t)}{\omega}),$ (3)
$b^{{\dagger}}(t)=\sqrt{\frac{\omega}{2}}(x(t)-\frac{ip(t)}{\omega}),$
$x(t)=\sqrt{\frac{1}{2\omega}}(b(t)+b^{{\dagger}}(t)),$
$p(t)=-i\sqrt{\frac{\omega}{2}}(b(t)-b^{{\dagger}}(t)),$
The boson number operator $b^{{\dagger}}(t)b(t)$ is $N$-by-$N$ Hermitian matrices with a highly degenerate spectrum and are completely degenerate initially (at t=0), reflecting the internal degrees of freedom. Also, we consider the case that the fermion number operator is always at an equilibrium steady state throughout the evolution to eliminate the potential effect of its fluctuation on the polaron shift. For electron-phonon interaction, we consider the Holstein mechanism
$H_{eb}(t)=g\omega c^{{\dagger}}(t)c(t)(b^{{\dagger}}(t)+b(t))=.g\sqrt{2\omega}\omega c^{{\dagger}}(t)c(t)x(t).$ (4)
@@ III.1 Polaronic shift
Different to the classical harmonic oscillator where the time-dependence of boson operator can be extracted as a phase factor, which is necessary for the Lang-Firsov transformation, the boson operator here cannot. Instead, we can separate the above boson operators (non-Hermitian) into the Hermitian part and non-Hermitian part, such that
$b(t)=b_{H}(t)+b_{nH}(t),$ (5)
$b^{{\dagger}}(t)=b^{{\dagger}}_{H}(t)+b^{{\dagger}}_{nH}(t),$
$x(t)=x_{1}(t)+x_{2}(t):=\sqrt{\frac{1}{2\omega}}(b_{H}(t)+b_{H}^{{\dagger}}(t))+\sqrt{\frac{1}{2\omega}}(b_{nH}(t)+b_{nH}^{{\dagger}}(t)),$
$p(t)=p_{1}(t)+p_{2}(t):=-i\sqrt{\frac{\omega}{2}}(b_{H}(t)-b_{H}^{{\dagger}}(t))-i\sqrt{\frac{\omega}{2}}(b_{nH}(t)-b_{nH}^{{\dagger}}(t)),$
where $[b_{H}(t),b^{{\dagger}}_{H}(t)]=0$ and $[b_{nH}(t),b^{{\dagger}}_{nH}(t)]=1$. Here the boson operators are still non-Hermitian and the position/momentum operators are still Hermitian. Note that ${\rm Tr}[x_{1}(t)^{2}]={\rm Tr}[x_{1}(t)]^{2}/N$ and ${\rm Tr}[x_{2}(t)^{2}]={\rm Tr}[x_{2}(t)]^{2}/N$, i.e., $x_{1}(t)$ and $x_{2}(t)$ are diagonally dominant matrices. We consider the case where the phase-space distribution remains, and as a product of the Gaussian Wigner distributions $x^{2}$ and $p^{2}$ which are inversely proportional, i.e., $\langle x^{2}\rangle=(4\langle p^{2}\rangle)^{-1}$, for the case of squeezed state at zero-temperature. This minimum uncertainty is guaranteed by the fact that both the total density matrix and the reduced one are all pure state during the evolution. Next we introduce the anti-Hermitian operator (such that $S^{{\dagger}}=-S$ and $e^{S}$ is unitary)
$S=c^{{\dagger}}(t)c(t)(b^{{\dagger}}(t)-b(t))=\frac{\sqrt{2}}{i\sqrt{\omega}}c^{{\dagger}}(t)c(t)p(t)$ (6)
and apply the Lang-Firsov transformation $\overline{O}=e^{S}Oe^{-S}$ to $H$, we obtain
$\overline{H}=\epsilon c^{{\dagger}}(t)c(t)+\omega b^{{\dagger}}(t)b(t)+\frac{{\bf I}}{2}+c^{{\dagger}}(t)c(t)\Sigma(t),$ (7)
where only the fermion number operator is invariant under the transformation. $\Sigma(t):=\Sigma_{eb}(t)+\Sigma_{b}(t)=g\omega e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}+(e^{S}H_{b}e^{-S}-H_{b})H_{b}^{-1}$ is the polaronic self-energy. This definition is consistent with Ref.[5], i.e., the transformed fermion-boson term and the differece between the fermion terms after and before transformation, which measures the shift of bosonic oscillator due to the presence of electron. Note that all the expectations in this article are in the single-electron basis state $c^{{\dagger}}(t)|0\rangle$. All the numerical simulations perfomed in this article base on a 2-by-2 electron Hilbert space and 4-by-4 boson Hilbert space, such that there is 8-dimensional combined basis ($N=8$), and we adopt the convenience of notation $c^{{\dagger}}\equiv c^{{\dagger}}_{2\times 2}\otimes{\bf I}_{4\times 4}$.
Due to the intrinsic unitarility we have ${\rm Tr}H_{b}(t=0)=N\omega$ and ${\rm Tr}H_{b}(t=\infty)=\omega$, where there is a smooth exponential decay with time. For $e^{S}x(t)e^{-S}$, diagonalization through the transformation of $e^{S}$ is only successful on $x_{2}(t)$,
$\frac{1}{g\omega N}{\rm Tr}[\Sigma_{eb}(0)]=\frac{1}{N}{\rm Tr}e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}=\frac{1}{N}{\rm Tr}e^{S}(b^{{\dagger}}_{H}(t)+b_{H}(t))e^{-S}+\frac{1}{N}{\rm Tr}e^{S}(b_{nH}^{{\dagger}}(t)+b_{nH}(t))e^{-S}$ (8)
$\sim\left\{\begin{array}[]{ll}2-{\rm ln}\omega,\ (\omega<1)\\ 2+(e^{-(\omega-1)}-1)\ (\omega>1)\end{array},\right.$
where the factor 2 originates from ${\rm Tr}[b_{H}(t=\infty)+b_{H}^{{\dagger}}(t=\infty)]=2$ (${\rm Tr}[b_{H}(t=0)+b_{H}^{{\dagger}}(t=0)]=2N$). The time and frequency dependence are important to seeking the minimal of polaronic shift. For $\omega\leq 1$ and at finite time,
$\frac{1}{g\omega N}{\rm Tr}[\Sigma_{eb}(t)]=\frac{1}{N}{\rm Tr}e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}=\frac{1}{N}{\rm Tr}e^{S}(b^{{\dagger}}_{H}(t)+b_{H}(t))e^{-S}+\frac{1}{N}{\rm Tr}e^{S}(b_{nH}^{{\dagger}}(t)+b_{nH}(t))e^{-S}$ (9)
$\sim\left\{\begin{array}[]{ll}(2-{\rm ln}\omega)\frac{\omega}{t},\ (small\ t)\\ (2-{\rm ln}\omega)e^{-t/\omega}\ (large\ t)\end{array}\right.+\left\{\begin{array}[]{ll}(2-{\rm ln}\omega)\frac{\omega}{t^{1/\omega}},\ (small\ t)\\ -\frac{(2-{\rm ln}\omega)e^{-t/\omega}+1}{\omega}\ (large\ t)\end{array},\right.$
where the early stage and latter stage follow the power-law decay and exponential decay, respectively. This is shown in Fig.2, in terms of the boson dispalcement. The expectation of $H_{b}$ in total density
@@
$\frac{1}{\omega}\langle H_{b}\rangle$$\sim\left\{\begin{array}[]{ll}\frac{1}{t},\ (small\ t)\\ e^{-t}\ (large\ t),\end{array}\right.$ (10)
as shown in Fig.1(a). Thus, the expectation exhibits a transition from power-law decay to exponential decay with time evolution. This signifies that the corresponding system density’s dynamics (Lindbladian) is gapless in the beginning and becomes gapped at long times. Note that during the whole process, the spectrum of the Lindbladian should be purely real (without the component of oscillation frequencies) due to the absence of non-local correlations and coherence (electron hopping) for $\omega\leq 1$.
Figure 1: (a) Expectation of boson operator on the single electron basis state. at long-time. (b)-(d) Expectations of position and momentum operators which follow the Gaussian Wigner distributions. (b) shows the case $\omega>1$ where the imaginary part of squared position and momentum operators emerges (this is also the regime where there is nonzero real part of expectation of momentum operator).
@@
Figure 2: Expectations of $x_{1}(t)$ and $x_{2}(t)$. The orange and green lines fit the early and latter stages using power-law and exponential decay, respectively.

### v4

# Non-Markovian to Markovian Transition in a Driven Single-Site Holstein Model

@@ Abstract
We investigate the non-adiabatic dynamics of the Holstein mechanism in a single-site model driven by a time-dependent perturbation. Unlike standard quenching protocols, we consider a unitary evolution where the driving field introduces a time-dependent displacement to the boson operators. The resulting dynamics reveal a dynamical quantum transition, evidenced by a distinct crossover from short-time non-Markovian behavior (characterized by power-law decay) to long-time Markovian-like behavior (exponential decay). This transition is consistently manifested in the polaronic shift, bosonic energy, and the population dynamics of the reduced density matrix. Furthermore, we analyze the emergence of effective non-unitary behavior. We demonstrate that while the global evolution remains unitary, the emergence of a degenerate ground state manifold at long times acts as an effective environment. Under a finite resolution threshold, the resulting fluctuations and mixing outweigh the coherent dynamics, leading to an effective unitary-to-nonunitary transition characterized by high mixedness and entanglement.
@@ I Introduction
The interaction between electrons and lattice vibrations (phonons) is a fundamental problem in condensed matter physics, leading to the formation of a quasi-particle known as the polaron, where an electron is "dressed" by a cloud of virtual phonons. The physical properties of these polarons, including their stability and phase transitions, are fundamentally influenced by whether the system is in an adiabatic or non-adiabatic regime [1]. The physics of polarons is broadly separated into two limits. In adiabatic limit where the electron’s Fermi velocity significantly exceeds the sound velocity, electron-phonon scattering is elastic, and interband transitions vanish, preventing electron energy loss [16, 17].
In this paper, we focus on the nonadiabatic regime, which becomes dominant under strong electron-phonon coupling. In this regime, a non-relativistic electron moving faster than the sound wave can interact with acoustic phonons, allowing it to lose energy via inelastic phenomena such as Cherenkov radiation[2]. This regime is crucial for understanding complex polaron formation, as strong non-adiabaticity can lead to significant modifications in the energy spectrum, such as the avoided crossing in the polaron band structure [7, 8]. Such avoided crossings and related strong-interaction phenomena have also been observed in ultracold Fermi atomic systems [9, 10].
We employ the single-site Holstein model as a canonical framework for investigating these nonadiabatic, strong-coupling polaron dynamics. While this model captures the essential physics of an electron’s self-trapping due to local interaction with bosons, our work introduces a novel perspective by studying the model under a time-dependent perturbation that drives the system’s unitary evolution. Specifically this time-dependent perturbation breaks the degeneracies of the composite electron-boson Hilbert space. It introduces a time-dependent Hermitian part (driving term) to the boson operators. This Hermitian contribution leads the ground state of the annihilation boson operator to deviate from the classical coherent Gaussian state, driving the system away from equilibrium.
The main goal of this work is to demonstrate that the resulting unitary dynamics reveal a quantum phase transition. We show that this transition manifests as a distinct crossover from non-Markovian behavior (characterized by power-law decay) at short times to Markovian behavior (exponential decay) at long times. This signature is numerically verified in key observables, including the polaronic shift, the bosonic energy, and the dynamics of the reduced density matrix. Unlike scenarios where decoherence arises from external bath effects or the orthogonality catastrophe in 1D systems [3, 4, 5], our single-site model provides a clean environment to study the system’s intrinsic non-Markovian to Markovian transition driven solely by the internal electron-phonon coupling under unitary evolution.
@@ II Model
We consider the single-site Holstein model in the Schrödinger picture. The total time-dependent Hamiltonian is given by
$H(t)=H_{0}+H_{p}(t),$ (1)
where the static part $H_{0}$ accounts for the electron, phonon, and their interaction
$H_{0}=\epsilon c^{\dagger}c+H_{b}+H_{eb}.$ (2)
Here, $c^{\dagger}$ ($c$) and $b^{\dagger}$ ($b$) are the time-independent creation (annihilation) operators for electrons and phonons, respectively.To investigate the non-equilibrium dynamics under non-adiabatic conditions, we introduce a time-dependent perturbation $H_{p}(t)$ that linearly drives the phonon field
$H_{p}(t)=f(t)(b^{\dagger}+b),$ (3)
where $f(t)$ is a time-dependent external field (e.g., a laser pulse). Here we introduce a time-dependent driving field $f(t)$ that couples to the lattice displacement. The perturbation is given by $H_{p}(t)=f(t)(b^{\dagger}+b)$. Since $(b^{\dagger}+b)\propto x$ corresponds to the Hermitian position quadrature, this term introduces a time-dependent Hermitian part to the boson dynamics, driving the system away from its equilibrium vacuum state. This term explicitly introduces the time-dependence into the system, breaking the time-translation symmetry and leading to the unitary evolution studied below. $\dot{b}(t)=i[H(t),b]=-i\omega b(t)-ig\omega c^{{\dagger}}c-if(t)$.
Due to the single-site consideration, the bare bandwidth is irrelevant here. Note that ${\rm Tr}[(c^{{\dagger}}c)^{2}]={\rm Tr}[c^{{\dagger}}c]^{2}/N$, precluding the Hubbrd-type interacion effect. We consider the total Hamiltonian in a N-by-N Hilbert space, and a time-dependent perturbation cause breaks the degeneracies of original composite Hilbert spaces. This guarantees the unitary evolution of $H(t)$, which is distinct from the Heisenberg picture. The bosonic term reads
$H_{b}=\frac{p^{2}}{2}+\frac{1}{2}\omega^{2}x^{2}=\omega(b^{{\dagger}}b+\frac{{\bf I}}{2}),$ (4)
where $b=\sqrt{\frac{\omega}{2}}(x+\frac{ip}{\omega}),\ b^{{\dagger}}=\sqrt{\frac{\omega}{2}}(x-\frac{ip}{\omega}),\ x=\sqrt{\frac{1}{2\omega}}(b+b^{{\dagger}}),\ p=-i\sqrt{\frac{\omega}{2}}(b-b^{{\dagger}})$. The boson number operator $b^{{\dagger}}b$ is NN-by-NN Hermitian matrices with a highly degenerate spectrum and are completely degenerate initially (at t=0), reflecting the internal degrees of freedom. Also, we consider the case that the fermion number operator is always at an equilibrium steady state throughout the evolution to eliminate the potential effect of its fluctuation on the polaron shift. For electron-phonon interaction, we consider the Holstein mechanism
$H_{eb}=g\omega c^{{\dagger}}c(b^{{\dagger}}+b)=g\sqrt{2\omega}\omega c^{{\dagger}}cx.$ (5)
@@ III.1 Polaronic shift
Different to the classical harmonic oscillator where the time-dependence of boson operator can be extracted as a phase factor, which is necessary for the Lang-Firsov transformation, the boson operator here cannot. Instead, we can separate the above boson operators (non-Hermitian) into the quantum fluctuation part and classical displacement part, such that
$b(t)=b(0)e^{-i\omega t}-i\int_{0}^{t}d\tau(g\omega c^{{\dagger}}c+f(\tau))e^{-i\omega(t-\tau)}=b_{0}(t)+\beta(t)-gc^{\dagger}c(1-e^{-i\omega t}),$ (6)
$b^{\dagger}(t)=b^{\dagger}(0)e^{i\omega t}+i\int_{0}^{t}d\tau(g\omega c^{{\dagger}}c+f(\tau))e^{i\omega(t-\tau)}=b_{0}^{\dagger}(t)+\beta^{*}(t)-gc^{\dagger}c(1-e^{i\omega t}),$
$x(t)=x_{1}(t)+x_{2}(t)+x_{3}(t)=\sqrt{\frac{1}{2\omega}}\left(b(0)e^{-i\omega t}+b^{\dagger}(0)e^{i\omega t}\right)+\sqrt{\frac{1}{2\omega}}\left(\beta(t)+\beta^{*}(t)\right)-gc^{\dagger}c\sqrt{\frac{2}{\omega}}(1-\cos(\omega t)),$
$p(t)=p_{1}(t)+p_{2}(t)+p_{3}(t)=-i\sqrt{\frac{\omega}{2}}\left(b(0)e^{-i\omega t}-b^{\dagger}(0)e^{i\omega t}\right)-i\sqrt{\frac{\omega}{2}}\left(\beta(t)-\beta^{*}(t)\right)-\sqrt{2\omega}gc^{{\dagger}}c{\rm sin}(\omega t),$
where we use $ig\omega c^{\dagger}c\int_{0}^{t}d\tau e^{-i\omega(t-\tau)}=-ig\omega c^{\dagger}c\left[\frac{e^{-i\omega(t-\tau)}}{i\omega}\right]_{0}^{t}=-gc^{\dagger}c(1-e^{-i\omega t})$, $ig\omega c^{\dagger}c\int_{0}^{t}d\tau e^{i\omega(t-\tau)}=-gc^{\dagger}c(1-e^{i\omega t})$. The commutation relation is satisfied: $\beta(t),\beta^{*}(t)]=[-gc^{\dagger}c(1-e^{-i\omega t}),-gc^{\dagger}c(1-e^{i\omega t})]=0$ and $b_{0}(t),b^{{\dagger}}_{0}(t)]=1$. Under the time-dependent perturbation $H_{p}(t)=f(t)(b^{\dagger}+b)$, the Heisenberg operators evolve as the sum of a free quantum evolution and a classical displacement and a polaron coupling-induced displacementy where $b_{0}(t)=b(0)e^{-i\omega t}$ is the free evolution operator and $\beta(t)=-i\int_{0}^{t}d\tau f(\tau)e^{-i\omega(t-\tau)}$ is the perturbation-induced displacement. The quantum Fluctuations ($x_{1}$ and $p_{1}$) originate from the free evolution $b(0)$, these terms represent the intrinsic quantum uncertainty. They satisfy ${\rm Tr}[x_{1}]=0$ but ${\rm Tr}[x_{1}^{2}]\neq 0$, preserving the commutation relations. The classical mean-fields ($x_{2}$, $p_{2}$ and $x_{3}$, $p_{3}$) arise from the external drive $f(t)$ and the electron-phonon coupling $g$, respectively. Within the single-electron subspace, these are proportional to the identity matrix (c-numbers). They satisfy the scalar trace relation ${\rm Tr}[O^{2}]=({\rm Tr}[O])^{2}/N$, representing well-defined classical trajectories in phase space.
Thus the position and momentum operators separate into quantum fluctuation parts ($x_{1}$ and $p_{1}$,), classical mean-field parts ($x_{2}$ and $p_{2}$) where $x_{2}(t)=\sqrt{\frac{2}{\omega}}\text{Re}[\beta(t)]$ and $p_{2}(t)=\sqrt{2\omega}\text{Im}[\beta(t)]$ are the time-dependent classical expectations driven by the external field, and the polaornic shift $x_{3}(t)$ and $p_{3}$. Note that all these components $x_{i},p_{i}$ ($i=1,2,3$) are Hermitian matrices. $x_{2}(t)$ is the unique term responsible for the relaxation dynamics (decay), whereas $x_{1}(t)$ and $x_{3}(t)$ represent persistent quantum fluctuations and coherent polaronic oscillations, respectively, which do not decay with time.
The classical parts $x_{2}(t)$ and $p_{2}(t)$ are proportional to the identity matrix in the boson subspace (c-numbers derived from the driving field), satisfying the relations ${\rm Tr}[x_{2}^{2}]=({\rm Tr}[x_{2}])^{2}/N$ and ${\rm Tr}[p_{2}^{2}]=({\rm Tr}[p_{2}])^{2}/N$. In contrast, the quantum parts $x_{1}(t)$ and $p_{1}(t)$ describe the fluctuations around the classical trajectory and do not satisfy this trace relation. We consider the case where the phase-space distribution remains a product of Gaussian Wigner distributions. The minimum uncertainty condition $\langle x_{1}^{2}\rangle=(4\langle p_{1}^{2}\rangle)^{-1}$ holds for the quantum fluctuation part ($\langle x_{1}^{2}\rangle=\frac{1}{2\omega}$, $\langle p_{1}^{2}\rangle=\frac{\omega}{2}$), consistent with the fact that the reduced density matrix describes a pure coherent state (displaced vacuum) throughout the unitary evolution.To decouple the electron-phonon interaction, we apply the standard Lang-Firsov (polaron) transformation, defined by the unitary operator $e^{S}$, with the anti-Hermitian generator
$S=gc^{\dagger}c(b^{\dagger}-b)=-ig\sqrt{2\omega}c^{\dagger}cp_{1}(0).$ (7)
Note that we define $S$ using the time-independent Schrödinger operators to establish a consistent polaron frame. Applying this transformation $\overline{O}=e^{S}Oe^{-S}$ to the total Hamiltonian $H(t)$, we obtain the transformed Hamiltonian
$\overline{H}_{0}=\epsilon c^{\dagger}c+\omega(b^{\dagger}-gc^{\dagger}c)(b-gc^{\dagger}c)+g\omega c^{\dagger}c(b^{\dagger}-gc^{\dagger}c+b-gc^{\dagger}c)=(\epsilon-g^{2}\omega)c^{\dagger}c+\omega b^{\dagger}b+\frac{\omega}{2},$ (8)
where we use $(c^{\dagger}c)^{2}=c^{\dagger}c$, $e^{S}c^{\dagger}ce^{-S}=c^{\dagger}c$, $e^{S}be^{-S}=b-[b,S]=b-gc^{\dagger}c$, $e^{S}b^{\dagger}e^{-S}=b^{\dagger}-gc^{\dagger}c$. The transformed driving term becomes
$\overline{H}_{p}(t)=f(t)[(b^{\dagger}-gc^{\dagger}c)+(b-gc^{\dagger}c)]=f(t)(b^{\dagger}+b)-2gf(t)c^{\dagger}c.$ (9)
For self-energy $\Sigma(t)$ arising from the external drive, it reads $\Sigma(t)=-g^{2}\omega-2gf(t)$. Thus the external perturbation acts as a time-varying potential on the dressed electron, modulating its binding energy $g^{2}\omega$.
Due to the unitary evolution generated by the time-dependent drive, the expectation value of the boson Hamiltonian does not correspond to a simple thermal decay but follows the dynamics of the displacement field. Next we analyze the polaronic shift. Applying the Lang-Firsov transformation, the effective electron-phonon coupling term (related to the self-energy) becomes
$\frac{1}{g\omega N}\langle\Sigma_{eb}(t)\rangle$ $=\frac{1}{N}\langle e^{S}(b^{\dagger}(t)+b(t))e^{-S}\rangle$ (10)
@@
$=\frac{1}{N}\langle\left[b^{\dagger}(0)e^{i\omega t}+b(0)e^{-i\omega t}\right]+[\beta^{*}(t)+\beta(t)]-gc^{\dagger}c-gc^{\dagger}c\rangle$
$=\frac{1}{N}\left((\beta^{*}(t)+\beta(t))-2g\right)$
where we use
$e^{S}b(t)e^{-S}$ $=e^{S}\left[b(0)e^{-i\omega t}+\beta(t)-gc^{\dagger}c+gc^{\dagger}ce^{-i\omega t}\right]e^{-S}$ (11)
$=\left(e^{S}b(0)e^{-S}\right)e^{-i\omega t}+\beta(t)-gc^{\dagger}c+gc^{\dagger}ce^{-i\omega t}$
$=(b(0)-gc^{\dagger}c)e^{-i\omega t}+\beta(t)-gc^{\dagger}c+gc^{\dagger}ce^{-i\omega t}$
$=b(0)e^{-i\omega t}+\beta(t)-gc^{\dagger}c,$
$e^{S}b^{{\dagger}}(t)e^{-S}$ $=b^{\dagger}(0)e^{i\omega t}+\beta^{*}(t)-gc^{\dagger}c,$
Note that $\langle b_{0}^{\dagger}(t)+b_{0}(t)\rangle=\langle b^{\dagger}(0)e^{i\omega t}+b(0)e^{-i\omega t}\rangle=0$, $e^{S}b_{0}(t)e^{-S}=e^{S}(b(0)e^{-i\omega t})e^{-S}=\left(e^{S}b(0)e^{-S}\right)e^{-i\omega t}=\left(b(0)-gc^{\dagger}c\right)e^{-i\omega t}$ and thus $e^{S}(b_{0}(t)+b_{0}^{{\dagger}}(t))e^{-S}=\left(b^{\dagger}(0)e^{i\omega t}+b(0)e^{-i\omega t}\right)-gc^{\dagger}c(e^{-i\omega t}+e^{i\omega t})=\left(b^{\dagger}(0)e^{i\omega t}+b(0)e^{-i\omega t}\right)-2gc^{\dagger}c\cos(\omega t)$. The transformed term involving the quantum fluctuations, $e^{S}(b_{0}^{\dagger}(t)+b_{0}(t))e^{-S}$, contains operators oscillating at frequency $\omega$. However, since the system remains in the vacuum state with respect to the fluctuation operators $b_{0}$, their expectation value vanishes, $\langle b_{0}^{\dagger}(t)+b_{0}(t)\rangle=0$. The non-trivial time-dependence of the polaronic shift $\Sigma(t)$ is entirely dominated by the classical displacement part $\beta(t)$, which yields the observed transition from power-law to exponential decay.
Note that the quantum fluctuation terms vanish under the expectation value in the vacuum state, i.e., $\langle b_{0}(t)\rangle=0$. The dynamics are thus strictly determined by the classical displacement $\beta(t)$. Under the driving protocol $f(t)$, the asymptotic behavior of the displacement is,
$\beta^{*}(t)+\beta(t)\sim\left\{\begin{array}[]{ll}(2-\ln\omega)\frac{\omega}{t},&(t\ll 1)\\ (2-\ln\omega)e^{-t/\omega},&(t\gg 1)\end{array}\right$. (12)
Fig.2 shows the time evolution of the classical component of the position operator $x_{2}(t)$. The solid lines represent the numerical results, while the orange and green dashed lines indicate the fitting for the early-stage power-law decay and late-stage exponential decay, respectively. This confirms that the macroscopic displacement dynamics are driven by the external perturbation.
By inverting the equation of motion $\dot{\beta}=-i\omega\beta-if(t)$, we can deduce that the observed transition in the displacement $\beta(t)$ directly reflects the spectral properties of the external perturbation $f(t)$. A driving field that transitions from a power-law decay ($f(t)\sim t^{-\alpha}$) to an exponential decay ($f(t)\sim e^{-\gamma t}$) will naturally induce the non-Markovian to Markovian crossover observed in Fig.2. From the relation $f(t)=i\dot{\beta}(t)-\omega\beta(t)$, the observed crossover in $\beta(t)$ from power-law ($f(t)\approx i\dot{\beta}(t)\approx i(-\alpha t^{-\alpha-1})-\omega t^{-\alpha}\approx-\omega t^{-\alpha}$) to exponential ($f(t)\approx i\dot{\beta}(t)\approx i(-\gamma e^{-\gamma t})-\omega e^{-\gamma t}=-(\omega+i\gamma)e^{-\gamma t}$) decay implies a similar transition in the temporal correlations of the perturbation $f(t)$. This confirms that our time-dependent perturbation effectively mimics a bath interaction that transitions from a noise regime of non-Markovian with long-range memory to a white noise regime (Markovian with short-range memory). This provides a dynamical origin for the observed phase transition in the polaron formation process.
This transition implies a crossover from power-law decay to exponential decay. The total boson energy expectation value behaves as:
$\frac{1}{\omega}\langle H_{b}(t)\rangle=\langle b^{\dagger}(t)b(t)\rangle+\frac{1}{2}=|\beta(t)|^{2}+\frac{1}{2}\sim\left\{\begin{array}[]{ll}\frac{1}{t},&(t\ll 1)\\ e^{-t},&(t\gg 1)\end{array}\right$. (13)
where $\langle b^{\dagger}(t)b(t)\rangle=\langle(b_{0}^{\dagger}(t)+\beta^{*}(t))(b_{0}(t)+\beta(t))\rangle=\langle b_{0}^{\dagger}b_{0}\rangle+\beta(t)\langle b_{0}^{\dagger}\rangle+\beta^{*}(t)\langle b_{0}\rangle+|\beta(t)|^{2}=|\beta(t)|^{2}$, as shown in Fig.1. Consequently, the particle number behaves as $\langle b^{\dagger}(t)b(t)\rangle\sim\frac{1}{t+1/N}$ for small $t$, and $\sim 1+e^{-t}$ for large $t$. This transition from power-law to exponential decay signifies that the system’s relaxation dynamics shift from a non-Markovian regime (memory effects) to a Markovian-like regime at long times, governed by the time-dependence of the external drive $f(t)$.
Figure 1: Expectation of boson operator on the single electron basis state. at long-time.
@@
Figure 2: Trace of the classical part of position operator $x_{2}(t)$. The orange and green lines fit the early and latter stages using power-law and exponential decay, respectively.